OpenStudy (anonymous):

What is the scale factor of a cube with a volume of 343m^3 to a cube with a volume of 5,832m^3? A. 324:49 B. 49:324 C: 18:7 D. 7:18 I think it is B.

OpenStudy (anonymous):

@jdoe0001 can you help me??

OpenStudy (jdoe0001):

\(\bf \large{ \cfrac{{\color{brown}{ volume}}}{{\color{blue}{ volume}}}=\cfrac{{\color{brown}{ side}}^3}{{\color{blue}{ side}}^3}\implies \cfrac{\sqrt[3]{{\color{brown}{ volume}}}}{\sqrt[3]{{\color{blue}{ volume}}}}=\cfrac{{\color{brown}{ side}}}{{\color{blue}{ side}}} \\ \quad \\ \quad \\ \cfrac{{\color{brown}{ 343}}}{{\color{blue}{ 5832}}}=\cfrac{{\color{brown}{ side}}^3}{{\color{blue}{ side}}^3}\implies \cfrac{\sqrt[3]{{\color{brown}{ 343}}}}{\sqrt[3]{{\color{blue}{ 5832}}}}=\cfrac{{\color{brown}{ side}}}{{\color{blue}{ side}}}\impliedby ratio }\)

OpenStudy (anonymous):

18^3/74^3

OpenStudy (jdoe0001):

is that what you got?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

well. the ratio will just be two single values

OpenStudy (jdoe0001):

what say does \(\large { \bf \cfrac{\sqrt[3]{{\color{brown}{ 343}}}}{\sqrt[3]{{\color{blue}{ 5832}}}} }\) give you?

OpenStudy (jdoe0001):

hint: use your calculator

OpenStudy (anonymous):

7/18

OpenStudy (anonymous):

It';s D

OpenStudy (jdoe0001):

then that's the ratio :)

OpenStudy (anonymous):

Can you help me with a few more??

OpenStudy (jdoe0001):

I'll be dashing in a few secs

OpenStudy (anonymous):

I only have 3

OpenStudy (anonymous):

Just those three